OpenStudy (anonymous):

what is the sum of the finite arithmetic series? (-10)+0+10+20+....+130

OpenStudy (amistre64):

n(first + last) ----------- 2

OpenStudy (amistre64):

in other words we can know 2 times the sum better, then divide it in half

OpenStudy (anonymous):

The answer choices are 900 1,000 950 955

OpenStudy (aum):

Can you figure out how many terms are in the sequence without using any formula?

OpenStudy (amistre64):

what does the formula determine as an answer then?

OpenStudy (amistre64):

icame up with at least 3 different number of terms in my head alone lol

OpenStudy (amistre64):

A = -10 + 0 + ...+130 +A = 130 + 120 + ... - 10 ------------------------ 2A = 120 + 120 + .... + 120, n times therefore A = 120n/2

OpenStudy (anonymous):

im not quite sure which one it is...

OpenStudy (aum):

10, 20 2 terms 10, 20, 30 3 terms 10, 20, 30, ...., 130 ? terms then include the first two terms.

OpenStudy (anonymous):

is it 900?

OpenStudy (aum):

Yes.